OpenStudy (anonymous):

What is the mean, median, and mode of 12,14,25,10,25,15,11

OpenStudy (anonymous):

well median is easy it is middle so line it up like this 10, 11, 12, 14, 15, 25, 25, and cros them off one by one so you get 14

OpenStudy (anonymous):

is mean 16?

OpenStudy (anonymous):

mode is like most minus less so you get 15 (i think haven't done this in a while)

OpenStudy (anonymous):

neither have I

OpenStudy (anonymous):

10+11+12+14+15+25+25=112/7=16 for the mean

OpenStudy (anonymous):

so yeah what skylar said for mean i remember it like teacher is mean for making us do this

OpenStudy (anonymous):

i believe the mode is the the number that ocurr the most

OpenStudy (anonymous):

which is 25

OpenStudy (anonymous):

The mode of a set of data is the value in the set that occurs most often. A set of data can be bimodal. It is also possible to have a set of data with no mode.